OpenStudy (anonymous):

A company's monthly profit, P, from a product is given by P = −x2 + 105x − 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550?

OpenStudy (alexwee123):

substitute $1550 for P then just factor out the equation o.0

OpenStudy (anonymous):

so 1550 = -x^2 + 105x - 1050? And then do i factor -x^2 + 105x - 1050?

OpenStudy (alexwee123):

yep :D

OpenStudy (alexwee123):

the you would get two zeros so pick the one with the least value and also it cannot be negative

OpenStudy (alexwee123):

then*

OpenStudy (anonymous):

wait what?! :o

OpenStudy (alexwee123):

wait hold on.... is the original equation P = −x2 + 105x − 1050?

OpenStudy (anonymous):

yes

OpenStudy (alexwee123):

then the equation you need to factor out is −x2 + 105x − 2100=0

OpenStudy (alexwee123):

ah well just solve it out with whatever method you want and tell me what you get

OpenStudy (anonymous):

but no factors of -2100 are the sum of 105? i think..

OpenStudy (anonymous):

do you know what the answer is? @alexwee123

OpenStudy (anonymous):

May I help you ??

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

You have given P = 1550: Plug this value in: \(1550 = -x^2 + 105x - 1050\) So: \(x^2 - 105x + 2600 = 0\) You have to solve this quadratic equation..

OpenStudy (anonymous):

Here a = 1, b = -105 and c = 2600 Find Discriminant here: \(D = b^2 - 4ac\) \(D = 11025 - 10400\) = ??

OpenStudy (anonymous):

So you will get like 625 after subtracting them right ??

OpenStudy (anonymous):

\[D = 625 \implies \sqrt{D} = 25\] Right ??

OpenStudy (anonymous):

yes correct!

OpenStudy (anonymous):

oh well i factored the equation x^2 - 105x + 2600 = 0 and got (x - 65) ( x - 40?

OpenStudy (anonymous):

(x - 65) ( x- 40)

OpenStudy (anonymous):

I did the method where you take the factors of 2600 that are the sum of -105

OpenStudy (anonymous):

Great well done..

OpenStudy (anonymous):

\[(x-65)(x-40) = 0\] So, x = 65 and x = 40 you will get..

OpenStudy (anonymous):

i'm confused :o

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

oh no okay i understand! did you make x - 65 = 0 and x - 40 = 0? Then you get x = 65 and x = 40 right?

OpenStudy (anonymous):

Yes I did the same..

OpenStudy (anonymous):

oh ok so then is the final answer 40?! :o

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

ok thank you :D